OpenStudy (anonymous):

summation of n=2 going to infinity of ((-1)^n sqt.root n)/ln(n). Determine whether the series is conditionally convergent, absolutely convergent, or divergent.

OpenStudy (anonymous):

\[\sum_{n=2}^{\infty}( (-1)^{n}\sqrt{n})/\ln (n)\]

OpenStudy (anonymous):

This is an alternating series, so you should use Leibniz criterion to determine the convergence: http://tutorial.math.lamar.edu/Classes/CalcII/AlternatingSeries.aspx